# Theory of Semi-Deterministic Quantum Dot Placement in Heteroepitaxy

### v2

@@ Abstract
Achieving deterministic placement of self-assembled quantum dots (QDs) during epitaxial growth is essential for the reliable and efficient fabrication of high-quality single-photon sources and solid-state cavity quantum electrodynamics (cQED) systems, yet it remains a significant challenge due to the inherent stochasticity of QD nucleation processes. In this work, we theoretically demonstrate (e.g., on a (001)-oriented substrate) that deterministic QD nucleation within a pristine growth region can be achieved by engineering the boundary geometry of that region. During epitaxial growth, adatoms initially move toward the boundary and promote the formation of primary QDs along the boundary, driven by curvature and diffusion anisotropy. The resulting primary QDs distribution will generate many-body interactions that dynamically reshape the chemical potential landscape for subsequently deposited adatoms, enabling the formation of secondary QDs within the pristine growth region. These findings provide a theoretical foundation for reliable patterning of high optical-quality QDs, with potential applications in next-generation quantum photonic devices.
I \@slowromancapi@. Introduction
@@
We propose an overdamped Langevin framework demonstrating that semi-deterministic QD placement is theoretically achievable through the combined effects of geometry-defined boundary profiles and the secondary fields generated by previously nucleated QDs along the boundary. These boundary-anchored QDs dynamically reshape the chemical potential landscape via repulsive interactions, significantly modulating adatom diffusion kinetics and altering adatom trajectories beyond the boundary region. This stochastic formalism provides a comprehensive model for semi-deterministic spatial organization of QDs, opening new avenues for constructing spatially ordered cQED systems for scalable quantum networks.
@@
II \@slowromancapii@. Adatom density and QD Nucleation in the Quasi-Stationary Limit
@@
In traditional heteroepitaxy, atoms are deposited onto a planar substrate with effectively boundaries at infinity, under homogeneous temperature and stress conditions. The spatial adatom density is typically approximated as translationally invariant. As a result, nucleation sites are uniformly distributed across the entire substrate. QD nucleation typically occurs via the Stranski–Krastanov (SK) growth mode [37, 38]. This mode is one of the three primary epitaxial growth modes observed in thin-film deposition and is particularly relevant in heteroepitaxy where two different materials are involved (e.g., InAs on GaAs) [39]. It describes a transition from two-dimensional (2D) layer-by-layer growth to three-dimensional (3D) island formation. When deposition begins, the adatoms initially form a thin, flat wetting layer on the substrate surface. This layer grows coherently, adopting the lattice constant of the underlying substrate despite a lattice mismatch. As the wetting layer thickens, strain energy accumulates due to lattice mismatch between the film and the substrate (e.g., InAs has a $\sim 7\%$ larger lattice constant than GaAs) [40]. After reaching a critical thickness, typically a few monolayers, the strain energy induces local instabilities that trigger the spontaneous formation of three-dimensional islands. Adatoms begin to aggregate at these energetically favorable sites and form nanometer-sized clusters, resulting in the formation of QDs. The density and size of these dots can be tuned based on the growth conditions.
@@
Adatoms exhibit anisotropic diffusion, and arsenic (As) adatoms tend to form surface dimers upon attachment, resulting in dimer row or valley structures. Indium (In) atoms diffuse with significantly different rates depending on whether their motion is tangential, which is slower and aligned along $\mathbf{\hat{y}}\equiv[1\bar{1}0]$, or orthogonal, which is faster and aligned along $\mathbf{\hat{x}}=[1{1}0]$, to the As dimer rows [41]. This directional anisotropy can lead to elongated quantum dots aligned along the fast diffusion axis $[1\bar{1}0]$ [43]. While in this work we focus on epitaxial growth on a (001)-oriented substrate, as reflected in the definition of crystallographic directions and the form of the diffusion tensor, the theoretical framework is in general applicable to substrates with arbitrary orientations, e.g. (111)-orientated surface [42].
@@
As suggested by several recent experimental studies, the introduction of boundary fields may arise from local strain gradients, such as those generated by buried stressors [25, 26] and capping layers [27, 28], or from thermal gradients across the substrate. As illustrated in Fig.1, we present several examples of boundary fields that can be parametrized either by arc length or by the shape of composite geometries. We classify boundary fields into three types: exterior, interior, and soft boundary fields. Both exterior and interior fields correspond to hard physical boundaries (for example, hard mask edges) that adatoms cannot physically penetrate during the epitaxy process. In contrast, a soft boundary field, such as one induced by buried stressors, is associated with partially forbidden regions where adatom migration is energetically unfavorable during growth, for example, in regions exhibiting thermal or strain gradients. Boundary field strengths, governed by in-plane compressive strain and local impurities, can alter the diffusion dynamics of adatoms during epitaxy, resulting in an increased likelihood of quantum dot nucleation within or near the defined boundary region. We refer to the exposed pristine surface (white-colored regions in Fig.1) as the bulk in this work.
@@
[…]
The adatom number conservation expressed above holds approximately under the assumption that the epitaxy rate is much slower than the timescale required for adatom diffusion to reach a quasi-steady state. To model the atom deposition process during epitaxy, we introduce adatoms uniformly across the exposed substrate area, with a deposition rate that is significantly smaller than the diffusion timescale. We refer to this regime as the quasi-stationary limit.
In the quasi-stationary limit, the evolution of the adatom density is governed by the coupled dynamics of the local adatom density $n(\mathbf{r})$ and the associated density flux $\mathbf{J}(\mathbf{r})$, described by the Boltzmann weight $P[n(\mathbf{r}),\mathbf{J}(\mathbf{r})]$,
$P[n,\mathbf{J}]$ $=\frac{1}{\mathcal{Z}}\exp\!\left[-\frac{1}{k_{B}T}\,H[n,\mathbf{J},B]\right],$ (2)
where $H[n(\mathbf{r}),\mathbf{J}(\mathbf{r}),B]$ is a general many-body correlation functional associated with the field configuration $[n,\mathbf{J}]$, and $\mathcal{Z}$ is the partition function, defined as the functional integral $\mathcal{Z}=\int\mathcal{D}n\,\mathcal{D}\mathbf{J}\;P[n,\mathbf{J}]$ over all admissible field configurations.
The QD nucleation probability at a site $\mathbf{r}$ depends on the ensemble average over all density and flux fields and takes the form,
$P_{\mathrm{QD}}(\mathbf{r})=\frac{1}{\mathcal{Z}}\int\mathcal{D}n\,\mathcal{D}\mathbf{J}\;P[n,\mathbf{J}]\,\lambda\big{(}n(\mathbf{r}),\mathbf{J}(\mathbf{r})\big{)},$ (3)
where $\lambda(n(\mathbf{r}),\mathbf{J}(\mathbf{r}))$ is the local nucleation rate function, modeled as,
$\lambda(n,\mathbf{J})=\lambda_{0}\cdot\Theta\big{(}n-n_{\mathrm{crit}}\big{)}\cdot\mathcal{R}(n,\mathbf{J}),$ (4)
with $n_{\mathrm{crit}}$ denoting the critical adatom density for QD nucleation. Here, $\Theta$ is the Heaviside step function, and $\mathcal{R}(n,\mathbf{J})$ is a modulation factor that may be modeled as a delta function $\delta(n-n_{\mathrm{crit}})$, or a smooth exponential dependence on $n$ and $\mathbf{J}$.
@@
III \@slowromancapiii@. Empirical Field, QD Nucleation, and Langevin-Based Monte Carlo Sampling
@@
where $m$ is the effective mass, $\boldsymbol{\Gamma}$ is the friction tensor (assumed diagonal) along the two principal crystallographic directions $[1\bar{1}0]$ and $[110]$, and $\boldsymbol{\eta}_{i}(t)$ is a Gaussian noise term (a distribution) with statistics $\langle\eta_{i}(t)\,\eta_{j}(t^{\prime})\rangle=\delta_{ij}\,\delta(t-t^{\prime})$.
The total microscopic potential energy $U_{\mathrm{sys}}$ includes both the boundary field $B$ and pairwise interactions $w$, defined as a functional of the empirical adatom density field over the entire exposed substrate area $\Omega$,
[…]
@@
where the diffusion tensor is given by $\mathbf{D}={\boldsymbol{\Gamma}^{-1}}$, and the functional derivative $\mu=\delta U_{\mathrm{sys}}/\delta\hat{n}(\mathbf{r},t)$ represents the single-adatom chemical potential, and its spatial gradient $-\nabla\mu$ corresponds to the microscopic force acting on the adatom ensemble. In this framework, the total force is modeled as the sum of a boundary-induced term and pairwise interaction contributions,
[…]
Although this expression is formally valid, the functional derivative is not rigorously defined due to the singular nature of $\hat{n}$, which consists of Dirac delta functions. In practice, force gradients are computed directly from the adatom-level Langevin dynamics. Thus, the SDE above should be interpreted as describing the stochastic evolution of an adatom ensemble, rather than as a coarse-grained partial differential equation (PDE).
To sample the density field $\hat{n}(\mathbf{r},t)$, we do not solve the above SDE directly. Instead, we simulate Langevin dynamics for an ensemble of $N$ adatoms and estimate the empirical field via a grid-based Monte Carlo technique. Specifically, we evolve the single-adatom Langevin equation given in Eq. 6,
@@
where $W_{i}(t)$ is a standard Wiener process. From the resulting trajectories, the empirical adatom density field is computed by counting particles on a spatial grid element $\mathrm{d}\Omega$, parametrized by the two-dimensional variable $\boldsymbol{l}$, centered at $\mathbf{r}_{0}$ and covering the exposed substrate region $\Omega$. This yields a time-dependent realization of the empirical field that satisfies the above SDE in expectation over ensemble averages,
@@
The above Langevin formulation, which generates the empirical adatom density field, must also incorporate microscopic nucleation processes that lead to QD formation during heteroepitaxy. To account for this, we define the local nucleation free energy $G(\hat{n},\mathbf{r}_{0})$ as the free energy cost to assemble a cluster of $\hat{n}$ adatoms within a grid cell of width $\sqrt{d\Omega}$, centered at $\mathbf{r}_{0}$. This nucleation barrier is computed by evaluating the system free energy $U_{\mathrm{sys}}[\hat{n}(\mathbf{r},t)]$ on a constrained configuration in which $n$ adatoms are localized near $\mathbf{r}_{0}$, and subtracting the reference free energy of the homogeneous background state $\hat{n}(\mathbf{r},0)\sim 1/{\Omega}$,
@@
The function $G(\hat{n},\mathbf{r}_{0})$ captures the energetic trade-off between inter-adatom interactions, entropic spreading, strain relaxation, and the influence of boundary fields near $\mathbf{r}_{0}$. Typically, $G(\hat{n},\mathbf{r}_{0})$ exhibits two local minima and one maximum with respect to $\hat{n}$, corresponding to metastable and stable nucleation states separated by a free energy barrier [46].
To model the nucleation process under thermal fluctuations, we introduce a localized seed potential $V_{\mathrm{seed}}$ at $\mathbf{r}_{0}$ whenever the empirical density exceeds a subcritical threshold $n_{\mathrm{seed}}$. This seed persists for a finite duration $\tau_{c}$, during which the system is allowed to either evolve toward the critical density $n_{\mathrm{crit}}$, corresponding to the second minimum of $G$, or decay back to the metastable uniform state. If the empirical density reaches $n_{\mathrm{crit}}$ within the time window $\tau_{c}$, the seed potential is irreversibly converted into a stable QD potential $V_{\mathrm{QD}}$, indicating successful traversal of the nucleation barrier. Otherwise, the potential is removed, corresponding to dissolution due to thermal fluctuations.
In general, the seeding potential $V_{\mathrm{seed}}$ is attractive, promoting local adatom accumulation and cluster formation. Once a QD reaches a characteristic size determined by the growth conditions, further adatom attachment becomes energetically unfavorable. Therefore, the QD potential $V_{\mathrm{QD}}$ is modeled as repulsive, reflecting the stabilization of the dot and the inhibition of further growth beyond its stationary size.
The above construction captures the stochastic nature of nucleation. This concept is formalized through a time-dependent nucleation potential centered at $\mathbf{r}_{0}$, evolving as,

### v4

# Semi-Deterministic Quantum Dot Placement in Heteroepitaxy

@@ Abstract
Achieving deterministic placement of self-assembled quantum dots (QDs) during epitaxial growth is essential for the reliable and efficient fabrication of high-quality single-photon sources and solid-state cavity quantum electrodynamics (cQED) systems, yet it remains a significant challenge due to the inherent stochasticity of QD nucleation processes. In this work, we theoretically and numerically demonstrate that deterministic QD nucleation within a pristine growth region, e.g., InAs on a (001)-oriented GaAs substrate, can be achieved by engineering the boundary geometry of that region. During epitaxial growth, adatoms initially move toward the boundary and promote the formation of primary QDs along the boundary, driven by curvature and diffusion anisotropy. The resulting primary QDs distribution will generate many-body interactions that dynamically reshape the chemical potential landscape for subsequently deposited adatoms, enabling the formation of secondary QDs within the pristine growth region. These findings provide a theoretical foundation for reliable patterning of high optical-quality QDs, with potential applications in next-generation quantum photonic devices.
I I. Introduction
@@
We propose an overdamped Langevin framework demonstrating that semi-deterministic QD placement is theoretically achievable through the combined effects of geometry-defined boundary profiles and the secondary fields generated by previously nucleated QDs along the boundary. These boundary-anchored QDs dynamically reshape the chemical potential landscape, significantly modifying diffusion kinetics of newly deposited adatoms. This results in secondary QD formation on the pristine region within the boundaries. This stochastic formalism provides a comprehensive model for semi-deterministic spatial organization of QDs, opening new avenues for constructing spatially ordered cQED systems for scalable quantum networks.
@@
II II. Adatom density and QD Nucleation in the Quasi-Stationary Limit
@@
In traditional heteroepitaxy, atoms are deposited onto a planar substrate with effectively boundaries at infinity, under homogeneous temperature and stress conditions. The spatial adatom density is typically approximated as translationally invariant. As a result, nucleation sites are uniformly distributed across the entire substrate. QD nucleation typically occurs via the Stranski–Krastanov (SK) growth mode [37, 38]. This mode is one of the three primary epitaxial growth modes observed in thin-film deposition and is particularly relevant in heteroepitaxy where two different materials are involved (e.g., InAs on GaAs) [39]. It describes a transition from two-dimensional (2D) layer-by-layer growth to three-dimensional (3D) island formation. When deposition begins, the adatoms initially form a thin, flat wetting layer on the substrate surface. This layer grows coherently, adopting the lattice constant of the underlying substrate despite a lattice mismatch. As the wetting layer thickens, strain energy accumulates due to lattice mismatch between the film and the substrate (e.g., InAs has a $\sim 7\%$ larger lattice constant than GaAs) [40]. After reaching a critical thickness, typically a few monolayers, local instabilities are induced to release strain energy elastically, triggering spontaneous formation of symmetric three-dimensional islands [41]. Adatoms begin to aggregate at these energetically favorable sites, forming nanometer-sized clusters (QDs) that are stable (self-limited in size) and exhibit various morphologies and transitions, such as the pyramid-to-dome transition, which have been observed and investigated both experimentally and theoretically [43, 44, 42]. As more adatoms are introduced into the system, coherent (elastic) QD growth becomes metastable, and incoherent (plastic) growth mechanisms, such as dislocation formation, ripening, and the adoption of elongated island shapes [45, 41], emerge. In this work, we focus on the coherent QD growth regime within the pristine surface where QDs are considered dislocation-free and adopt symmetric shapes on the wetting layer.
@@
While the proposed theoretical framework is in general applicable to any substrates with arbitrary orientations, specifically, we address specifically the standard SK growth mode of InAs on the GaAs(001)-$\beta_{2}(2\times 4)$ surface [46] : arsenic (As) adatoms tend to form surface dimers upon attachment, resulting in dimer row or valley structures. Indium (In) atoms diffuse with significantly different rates depending on whether their motion is tangential, which is slower and aligned along $\mathbf{\hat{y}}\equiv[1{1}0]$, or orthogonal, which is faster and aligned along $\mathbf{\hat{x}}\equiv[1\bar{1}0]$, to the As dimer rows [47]. This biaxial anisotropy is governed by a rank-2 diffusion tensor, which results in elongated quantum dots aligned along the fast diffusion axis $1\bar{1}0] [$48]. In this work, we assume that the diffusion tensor is dominated by its diagonal components, although, in general, non-zero off-diagonal terms can contribute to surface diffusion, depending on the substrate crystallographic orientation, adatom species, and temperature.
@@
As suggested by several recent experimental studies, the introduction of boundary fields may arise from local strain gradients, such as those generated by buried stressors [25, 26] and capping layers [27, 28], or from thermal gradients across the substrate. As illustrated in Fig.1, we present several examples of boundary fields that can be parametrized either by arc length or by the shape of composite geometries. We classify boundary fields into three types: exterior, interior, and soft boundary fields. Both exterior and interior fields correspond to hard physical boundaries (for example, hard mask edges) that adatoms cannot physically penetrate during the epitaxy process. In contrast, a soft boundary field, such as one induced by buried stressors, is associated with partially forbidden regions where adatom migration is energetically unfavorable during growth, for example, in regions exhibiting thermal or strain gradients. Boundary field strengths, governed by in-plane compressive strain and local impurities, can alter the diffusion dynamics of adatoms during epitaxy, resulting in an increased likelihood of quantum dot nucleation within or near the defined boundary region. We refer to the exposed pristine surface (white-colored regions in Fig.1) as the bulk.
@@
where $\mathbf{J}(\mathbf{r},n,B,t)$ is the local adatom current density, $\mathbf{r}$ is the spatial coordinate, $t$ is time, and $B=\int du\,B[\mathbf{r}-\mathbf{r}_{\mathrm{bd}}(u)]$ represents the spatially varying boundary field strength at $\mathbf{r}$. The boundary field is parameterized by $u$, the arc-length variable along the boundary, and $\mathbf{r}_{\mathrm{bd}}(u)$ denotes the unit-speed spatial parametrization of the boundary at $u$. We note that, in general, the boundary field is a tensor that captures structural coupling within the boundary contour. Without loss of generality of our method, we assume that $B[\mathbf{r}-\mathbf{r}_{\mathrm{bd}}(u)]$ is a scalar field.
The current density (flux) $\mathbf{J}(\mathbf{r},n,B,t)$ can be decomposed into two contributions: the lateral current density $\mathbf{J}_{\mathrm{lateral}}$ and the vertical current density $\mathbf{J}_{\mathrm{vertical}}$, representing in-plane adatom diffusion and out-of-plane deposition or desorption, respectively. In the quasi-stationary limit, the epitaxy rate at a given site $\mathbf{r}$ is much smaller than the lateral diffusion rate, such that,
$|\mathbf{J}_{\mathrm{vertical}}|\sim\lambda n+J_{\mathrm{dep}}\left(\sum_{m=-\infty}^{\infty}\delta(t-m\,t_{\mathrm{dep}})\right)\ll|\mathbf{J}_{\mathrm{lateral}}|,$ (2)
where $\lambda$ is the constant desorption rate, $t_{\mathrm{dep}}$ is the period between subsequent depositions, and $J_{\mathrm{dep}}$ is the deposition flux magnitude. This leads to a separation of fast and slow timescales: the lateral adatom density equilibrates rapidly, while the source and sink terms in the continuity equation associated with deposition or desorption act only as slow-varying perturbations.
In the quasi-stationary limit, where $\mathbf{J}(\mathbf{r},n,B,t)\approx\mathbf{J}_{\mathrm{lateral}}$, the lateral evolution of the adatom density is governed by the coupled dynamics of the local adatom density $n(\mathbf{r})$ and the associated lateral current density $\mathbf{J}$, which can be described by a quasi-equilibrium Boltzmann distribution,
$P[n,\mathbf{J}]=\frac{1}{\mathcal{Z}}\exp\!\Big[-\frac{1}{k_{B}T}\,H[n,\mathbf{J},B]\Big],$ (3)
where $H[n(\mathbf{r}),\mathbf{J}(\mathbf{r}),B]$ is a general many-body correlation functional associated with the configuration $n,\mathbf{J}]$, and $\mathcal{Z}$ is the partition function $\mathcal{Z}=\int\mathcal{D}n\mathcal{D}\mathbf{J}\exp[-H/{k_{B}T}]$, integrated over all admissible field configurations.
The QD nucleation probability at a site $\mathbf{r}$ depends on the ensemble average over all density and current fields and takes the form,
$P_{\mathrm{QD}}(\mathbf{r})=\frac{1}{\mathcal{Z}}\int\mathcal{D}n\,\mathcal{D}\mathbf{J}\;P[n,\mathbf{J}]\,\cdot\nu\big(n(\mathbf{r}),\mathbf{J}(\mathbf{r})\big),$ (4)
where $\nu(n(\mathbf{r}),\mathbf{J}(\mathbf{r}))$ is the local nucleation rate function, modeled as,
$\nu(n,\mathbf{J})=\nu_{0}\cdot\Theta\big(n-n_{\mathrm{seed}}\big)\cdot\mathcal{R}(n,\mathbf{J}),$ (5)
with ${n}_{\mathrm{seed}}$ denoting the lowest adatom density (subcritical density) that initiates QD nucleation process at a given site. $\Theta$ is the Heaviside step function, and $\mathcal{R}(n,\mathbf{J})$ is a modulation factor that may be modeled as a delta function $\delta(n-{n}_{\mathrm{seed}})$, or a smooth exponential dependence on $n$ and $\mathbf{J}$.
@@
III III. Empirical Field, QD Nucleation, and Langevin-Based Monte Carlo Sampling
@@
Figure 2: The free energy landscapes with four regimes (shaded with different colors) of quantum dot formation, detailed in the main text. The formation self-limited QDs exhibits two local minima representing a dissolved state and self-limited QD state, separated by a free energy barrier, peaked (local maximum) at the density $\hat{n}_{\mathrm{LaMer}}$. As discussed in Eq. 14, the LaMer free energy $\Delta G_{\mathrm{LaMer}}$ (purple curve) represents, as a function of cluster density $\hat{n}$, the balance between the free energy contributions from interfacial elasticity (cyan curve) and the formation of the cluster surface (red curve). Four regimes (A-D) are separated by three characteristic densities, subcritical density $\hat{n}_{\mathrm{seed}}$, supercritical density $\hat{n}_{\mathrm{supercritical}}$, and self-limited density $\hat{n}_{\mathrm{self-limited}}$, associated with the formation of temporary clusters, clusters lead to coherent QD growth, and self-limited (stable) QDs. In regime D (shaded in blue), the net free energy $\Delta G$ given in Eq. 15 may exhibit a convex behavior around the self-limited (stable) QDs. $\hat{n}_{\mathrm{supercritical}}$ and $\hat{n}_{\mathrm{self-limited}}$ are referred as a single critical density $\hat{n}_{\mathrm{crit}}$.
@@
where $m$ is the effective mass, $\boldsymbol{\Gamma}$ is the friction tensor (assumed diagonal) along the two principal crystallographic directions $1\bar{1}0]$ and $[110]$, and $\boldsymbol{\eta}_{i}(t)$ is a Gaussian noise term (a distribution) with statistics $\langle\eta_{i}(t)\,\eta_{j}(t^{\prime})\rangle=\delta_{ij}\,\delta(t-t^{\prime})$.
The total microscopic potential energy $U_{\mathrm{sys}}$ includes both the boundary field $B$ and repulsive pairwise adatom interaction $w$, defined as a functional of the empirical adatom density field over the entire exposed substrate area $\Omega$,
$U_{\mathrm{sys}}[\hat{n}$] $=\int du\int_{\Omega}B[\mathbf{r}-\mathbf{r}_{\mathrm{bd}}(u)]\,\hat{n}(\mathbf{r},t)\,d\mathbf{r}$ (8)
$\quad+\frac{1}{2}\iint_{\Omega}\hat{n}(\mathbf{r},t)\,w(\mathbf{r}-\mathbf{r}^{\prime})\,\hat{n}(\mathbf{r}^{\prime},t)\,d\mathbf{r}\,d\mathbf{r}^{\prime}.$
@@
where the diffusion tensor is given by $\mathbf{D}={\boldsymbol{\Gamma}^{-1}}$, and the functional derivative $\mu=\delta U_{\mathrm{sys}}/\delta\hat{n}(\mathbf{r},t)$ represents the single-adatom chemical potential landscape, as the sum of a boundary-induced term and pairwise interaction contributions,
$\mu(\mathbf{r},t)=$ $\int du\,B\big[\mathbf{r}-\mathbf{r}_{\mathrm{bd}}(u)\big]+\int_{\Omega}w\big(\mathbf{r}-\mathbf{r}^{\prime}\big)\,\hat{n}(\mathbf{r}^{\prime},t)\,d\mathbf{r}^{\prime}.$ (10)
The spatial gradient of the above chemical potential landscape $\mathbf{F}(\mathbf{r},t)=-\nabla\mu(\mathbf{r},t)$ corresponds to the microscopic force acting on the adatom ensemble. While the functional derivative $\mu=\delta U_{\mathrm{sys}}/\delta\hat{n}(\mathbf{r},t)$ is not rigorously defined due to the singular nature of $\hat{n}$, which consists of Dirac delta functions, in practice, the values and gradients of $\mu(\mathbf{r},t)$ are computed directly from the adatom-level Langevin dynamics. Thus, Eq. 9 should be interpreted as describing the stochastic evolution of an adatom ensemble, rather than as a coarse-grained partial differential equation (PDE).
To sample the density field $\hat{n}(\mathbf{r},t)$, we do not solve Eq. 9 directly. Instead, we simulate Langevin dynamics for an ensemble of $N$ adatoms and estimate the empirical field via a grid-based Monte Carlo technique. Specifically, we evolve the single-adatom Langevin equation given in Eq. 7,
@@
where $W_{i}(t)$ is a standard Wiener process. From the resulting trajectories, the empirical adatom density field is computed by counting particles on a spatial grid element $\mathrm{d}\Omega$, parametrized by the two-dimensional variable $\boldsymbol{l}$, centered at $\mathbf{r}_{0}$ and covering the exposed substrate region $\Omega$. This yields a time-dependent realization of the empirical field that satisfies the above SDE over ensemble averages,
@@
The above Langevin formulation, which generates the empirical adatom density field, must also incorporate microscopic nucleation processes that lead to QD formation during heteroepitaxy. To account for this, we define the local nucleation free energy $G(\hat{n},\mathbf{r}_{0})$ as the free energy cost to assemble a cluster of $\hat{n}$ adatoms within a grid cell of width $\sqrt{d\Omega}$, centered at $\mathbf{r}_{0}$. This nucleation barrier is computed by evaluating the system energy $U_{\mathrm{sys}}[\hat{n}(\mathbf{r},t)]$ on a constrained configuration in which $\hat{n}$ adatoms are localized near $\mathbf{r}_{0}$, and subtracting the reference energy of the homogeneous background state $\hat{n}(\mathbf{r},0)\sim 1/{\Omega}$,
@@
The function $G(\hat{n},\mathbf{r}_{0})$ captures the energetic trade-off between inter-adatom interactions, strain relaxation, and the influence of boundary fields near $\mathbf{r}_{0}$. We denote the energy $U_{\mathrm{sys}}[\hat{n}(\mathbf{r}_{0},t)]$ as the free energy, as the entropic contribution (compared to the strain and surface energies) is negligible under typical growth conditions [45].
The apparent complexity of the above free energy functional can be simplified within the coherent SK growth mode by adopting a classical LaMer nucleation model [52, 53], which captures the essential competing mechanisms during coherent QD growth, and QDs are considered dislocation-free and exhibit symmetric shapes. Since the QD volume is directly proportional to the number of adatoms $\hat{n}$, for QD formation at a site $\mathbf{r}_{0}$ with an approximate nucleus diameter $d_{\mathrm{QD}}\sim\hat{n}^{1/3}$, the following model describes the competition between elastic strain energy relief and surface free energy, with magnitudes $\mathcal{E}_{\mathrm{el}}$ and $\gamma_{\mathrm{surf}}$, respectively,
$\Delta G_{\mathrm{LaMer}}(d_{\mathrm{QD}})=-\epsilon_{0}^{2}\mathcal{E}_{\mathrm{el}}[d_{\mathrm{QD}}]^{3}+\gamma_{\mathrm{surf}}[d_{\mathrm{QD}}]^{2},$ (14)
where $\epsilon_{0}=(a_{\mathrm{film}}-a_{\mathrm{sub}})/a_{\mathrm{sub}}$ is the lattice mismatch between the film ($a_{\mathrm{film}}$) and the substrate ($a_{\mathrm{sub}})$ lattice constants. The first term corresponds to the elastic strain energy relief with respect to the wetting layer, and the second term captures the energetic penalty of generating QD surfaces [53]. Motivated by the experimental observation of self-limited QDs [44], we introduce a phenomenological free energy term $\Delta G_{\mathrm{self-limited}}>0$, which limits coherent QD growth and accounts for the emergence of a second local minimum in the free energy landscape,
$\Delta G(d_{\mathrm{QD}},\mathbf{r}_{0})$ $\sim\Delta G_{\mathrm{LaMer}}(d_{\mathrm{QD}},\mathbf{r}_{0})+\Delta G_{\mathrm{self-limited}}(d_{\mathrm{QD}},\mathbf{r}_{0}).$ (15)
While attempts have been made in obtaining the approximated form of this term [54], the exact mechanisms underlying self-limited QD sizes remain under debate (e.g., the competition between kinetic and thermodynamic processes) and are beyond the scope of this work [55, 56]. Nevertheless, regardless of the detailed form of this free energy term, its existence, along with the engineered boundary field, plays a crucial role in manipulating the empirical density field of subsequently deposited adatoms, enabling semi-deterministic QD placement on the bulk surface.
As illustrated in Fig. 2, we present four regimes of the free energy $\Delta G$ that capture essential aspects of QD formation:
  • Regime A (kinetic-driven seeding): Adatoms are continuously deposited onto the bulk surface, where they stochastically migrate and contribute to the formation of surface layers on top of the substrate. Upon approaching the critical thickness, adatoms form temporary seeding clusters with subcritical density $\hat{n}_{\mathrm{seed}}$ as the starting seed of QD growth.
  • Regime B (seeding stabilization): Upon reaching the subcritical bulk volume, further adatom deposition leads to the formation of transient seeding clusters that are energetically favorable. These clusters may increase in local density $\hat{n}$ as they grow elastically. However, they remain susceptible to dissolution through thermal fluctuations or Ostwald ripening, in which smaller clusters are absorbed by larger ones to reduce the system’s total surface energy, until they eventually exceed a supercritical density $\hat{n}_{\mathrm{supercritical}}$ [55].
  • Regime C (coherent QD growth): Once clusters exceed the supercritical size, they grow continuously until reaching a self-limited density $\hat{n}_{\mathrm{self-limited}}$, at which point they become stable QDs. We note that the lattice mismatch must be sufficiently large to support stable QDs [45]. Although the stabilization mechanisms are still in debate, the second minimum of the free energy, $\Delta G_{\mathrm{LaMer}}+\Delta G_{\mathrm{self-limited}}$, accounts for the existence of self-limited QDs. In contrast, the classical LaMer nucleation model, $\Delta G_{\mathrm{LaMer}}$, predicts unbounded QD growth corresponding to a global minimum. Nevertheless, both the existence of self-limited QDs and their influence on adatoms and neighboring QDs have been investigated extensively, both theoretically and experimentally.
  • Regime D (self-limited QD formation): QDs reach a stationary size, and the addition of further adatoms is no longer energetically favorable (i.e., the QDs become metastable toward plastic growth), generating an effective repulsive force on nearby adatoms [56]. These QDs act as sources of strain fields that elastically distort the wetting layer and substrate [57]. As a consequence, a uniform spatial distribution of QDs emerges, as observed both experimentally and theoretically via kinetic Monte Carlo simulations [58, 59].
While obtaining all microscopic parameters associated with the above four regimes is challenging, we can obtain valuable insights on how to manipulate QDs by considering the following simplifications together with the aforementioned Langevin dynamics, Eq. 11. The above four regimes can be formalized through a time-dependent nucleation potential centered at $\mathbf{r}_{0}$, evolving as,